OpenStudy (anonymous):

evaluate the integral x^3+x/x dx

hartnn (hartnn):

separate out the numerator (x^3+x)/x = x^3/x + x/x

hartnn (hartnn):

then you can use x^n formula for both the terms

hartnn (hartnn):

or is it x^3 +(x/x) ?

OpenStudy (anonymous):

the notation isn't great

OpenStudy (anonymous):

probably (x^3+x)/x since the other way would probably have x/x =1

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x ^{3}+x }{ x }dx=\int\limits_{}^{}x ^{2}+1dx=\frac{ x ^{3} }{ 3 }+x+C\]